OpenStudy (anonymous):

A waiter earned $31 in tips in an hour. He was paid in $1 bills and $5 bills. Use a model to determine how many $5 bills he received if he received a total of 15 bills. 1 2 3 4

OpenStudy (goldphenoix):

15/5= ?

OpenStudy (anonymous):

Let x= number of 1s. Let y= number of 5s x + y = 15 and 1x + 5y = 31 Solve these simultaneos equations. x + y = 15 gives us x = 15 - y Substitute in the other: (15 - y) + 5y = 31 4y = 16 y = 4 4 five dollar bills

OpenStudy (anonymous):

15/5=3

OpenStudy (goldphenoix):

Bangkok response sounds more legit. >_> Let me think.

OpenStudy (anonymous):

Bangkok has the right answer.

OpenStudy (anonymous):

@Elysium